# Online Learning of Whittle Indices for Restless Bandits with Non-Stationary Transition Kernels

### v1

@@ Appendix B Proof of Theorem 1
is a non-increasing function of the window size $w$. This is because $C^{(n)}_{t,w}(s,a)$ is a non-decreasing function of $w$. This conludes the proof.
Appendix C Proof of Theorem 2
To prove Theorem 2, we first bound
$g_{t,n}(s,a,w)$ $=\mathbb{E}_{P_{n,t},\pi_{n,t}}\left[\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w}(s,a)}}\right]\overset{(a)}{\leq}\frac{H}{\sqrt{w_{n}\mathrm{p}_{min}}},$ (32)
where $(a)$ holds because there exists a positive probability $\mathrm{p}_{min}$ to visit every $(s,a)\in\mathcal{Z}_{2}$ at least once in every episode $t\in[T]$ for all arms $n\in[N]$, so $C^{(n)}_{t,w}(s,a)\geq w_{n}\mathrm{p}_{min}$ and it always satisfies that $\alpha_{t}^{(n)}(s,a)\geq H$.
Then, we can have
$\sum_{t=1}^{T}\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}$
$\leq$ $\sum_{t=1}^{T}\bigg{(}\frac{H}{\sqrt{w_{n}\mathrm{p_{min}}}}+w_{n}\epsilon_{n}\bigg{)}.$ (33)
After that, by substituting $\epsilon_{n}=T^{-k}$ and $w_{n}=T^{q}$, we get
$\sum_{t=1}^{T}\bigg{(}\frac{H}{\sqrt{w_{n}\mathrm{p_{min}}}}+w_{n}\epsilon_{n}\bigg{)}=HT^{1-q/2}+HT^{q+1-k}.$ (34)
Further, by substituting $q=\min\{2k/3,1\}$, we have
$HT^{1-q/2}+HT^{q+1-k}\leq HT^{r},$ (35)
where $r=\max\{(3-k)/3,1/2\}$. Hence,
$\sum_{t=1}^{T}\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}\leq HT^{r}.$ (36)
By substituting the above upper bound in Theorem 1, we can get Theorem 2.
Appendix D Experimental Setup
@@
Firstly, we discuss how we compute the regret in our numerical studies. Given a policy, we evaluate the regret of the policy at each episode $t$ by subtracting the total discounted sum reward of all arms obtained by the policy from the total discounted sum reward of all arms received by a Whittle index policy with access to the true transition probabilities. In our simulation, we have used MATLAB.
In all of our experiments, a discount factor $\gamma=0.99$ is considered. Moreover, we consider $2$ class of arms. In each class, the arms share same transition kernels and same reward function in each episode, which is a common setup in the RMAB literature [5, 23, 20]
Our policy can be computationally intensive due to the kernel maximization problem (15). In the next section, along with more experimental setups, we discuss some approaches to reduce the complexity for One Dimensional Bandit and Wireless Scheduling simulations.
D.1 One Dimensional Bandit
In One Dimensional Bandit applications, we can show that $Q_{n,\lambda,t}(s,0)$ is a decreasing function of $p_{n}(t)$, and $Q_{n,\lambda,t}(s,1)$ is an increasing function of $q_{n}(t)$. This is because the reward $r(s,a)$ is an increasing function of state $s$, which results in $V_{n,\lambda,t}(s)$ being an increasing function of state $s$. Using this structure, we can get closed-form solution of the maximization problem (15).
Recall that, in the numerical study of One Dimensional Bandit, for half of the $N$ arms, we consider the following kernel dynamics which are unknown and must be learned from data - (i) $p_{n}(t)$ changes to $\min\{p_{n}(t-1)+\epsilon_{n},1\}$ with probability $0.7$, or it changes to $\max\{p_{n}(t-1)-\epsilon_{n},0\}$ with probability 0.3; the initial value of $p_{n}(t)=0.5$ is used with $\epsilon_{n}=0.05$ and (ii) $q_{n}(t)=0.5$ is unknown but stationary. For the other half, probabilities $p_{n}(t)=0.5$ and $q_{n}(t)=1$ are known to the decision maker and do not change over episode $t$.
D.2 Wireless Scheduling
@@
For our wireless scheduling problem, the value function $V_{n,\lambda,t}(s)$ is an increasing function of $q_{n}(t)$ because the reward function $r(s,a)$ is a decreasing function of the AoI state $s$. This structure allows us to obtain a closed-form solution for the maximization problem (15). Furthermore, a closed-form expression for the Whittle index exists [4]. Though the closed-form expression is obtained for average cost problem, we can use the Whittle index expression from [4] when the discount factor $\gamma\rightarrow 1$. In our experiment, as we have used $\gamma=0.99$, the approximation does not affect the performance of our policy. Consequently, we can determine the Whittle index for our problem using a low-complexity method that exploits these closed-form expressions.
Recall that in the numerical study of wireless scheduling, we assume that $q_{n}(t)$ is unknown and non-stationary for half of the sources, whereas it is is unknown but stationary for the remaining half. For non-stationary arms, the variance of signal values $\sigma_{n}^{2}=0.9$ is used and the probability of successful transmission $q_{n}(t)$ changes to $\min\{q_{n}(t-1)+\epsilon_{n},1\}$ with probability $0.7$, or it changes to $\max\{q_{n}(t-1)-\epsilon_{n},0\}$ with probability 0.3; the initial value of $q_{n}(t)=0.1$ is used with $\epsilon_{n}=0.05$. For the other half, $q_{n}(t)=1$ is unknown but stationary.

### v2

@@ A.2 Proof of Theorem 1
is a non-increasing function of the window size $w$. This is because $C^{(n)}_{t,w}(s,a)$ is a non-decreasing function of $w$.
Now, in the above analysis $V_{n}$ is the optimal choice of drift values. Specifically, the optimal choice of $V_{n}$ from the set of drift values satisfies: $\max_{(s,a)\in\mathcal{S}\times A}\sum_{s^{\prime}\in\mathcal{S}}\bigg|P_{n,t}(s^{\prime}|s,a)-P_{n,t-1}(s^{\prime}|s,a)\bigg|\leq V_{n}/T\leq\max_{(s,a)\in\mathcal{S}\times A}\sum_{s^{\prime}\in\mathcal{S}}\bigg|P_{n,t}(s^{\prime}|s,a)-P_{n,t-1}(s^{\prime}|s,a)\bigg|+2/J_{n}T.$ Consequently, the optimal upper bound of total variation budget $V_{n,T}$ satisfies
$\tilde{V}_{n}\leq V_{n}\leq\tilde{V}_{n}+2T/J_{n}.$
Then, the upper bound becomes
$\sum_{t=1}^{T}O\bigg(\sum_{n=1}^{N}2|\mathcal{S}|G_{t,n}(w_{n})\bigg)+\sum_{n=1}^{N}O\bigg(w_{n}(\tilde{V}_{n}+2T/J_{n})H\bigg).$
A.3 Proof of Theorem 2
Lets denote the probability to visit every $(s,a)\in\mathcal{S}\times\mathcal{A}$ at least once in an episode for all arms $n\in[N]$ by $P_{\mathrm{min}}$. According to the condition in Theorem 2, $P_{\mathrm{min}}>0$.
Now, to prove Theorem 2, we bound
$g_{t,n}(s,a,w)$ $=\mathbb{E}_{P_{n,t},\pi_{n,t}}\left[\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w}(s,a)}}\right]\leq H\mathbb{E}_{P_{n,t},\pi_{n,t}}\left[\frac{1}{\sqrt{C^{(n)}_{t,w}(s,a)}}\right],$ (28)
where the number of visit $\alpha_{t}^{(n)}(s,a)$ in one episode is upper bounded by the time horizon $H$.
Let $\mathbb{E}[C^{(n)}_{t,w}(s,a)]=\mu$. Then, $\mu\geq 1$ because by definition, $C^{(n)}_{t,w}(s,a):=\max\bigg\{\sum_{s^{\prime}\in\mathcal{S}}C_{t,w}^{(n)}(s^{\prime},a,s),1\bigg\}.$ Moreover, $\mu\geq w_{n}P_{\mathrm{min}}$.
Now, we have
$\mathbb{E}\left[\frac{1}{\sqrt{C^{(n)}_{t,w}(s,a)}}\right]$ $=\mathbb{E}\left[\frac{1}{\sqrt{C^{(n)}_{t,w}(s,a)}}\bigg|C^{(n)}_{t,w}(s,a)<\frac{\mu}{2}\right]P\left(C^{(n)}_{t,w}(s,a)<\frac{\mu}{2}\right)$
$+\mathbb{E}\left[\frac{1}{\sqrt{C^{(n)}_{t,w}(s,a)}}\bigg|C^{(n)}_{t,w}(s,a)\geq\frac{\mu}{2}\right]P\left(C^{(n)}_{t,w}(s,a)\geq\frac{\mu}{2}\right)$ (29)
If $C^{(n)}_{t,w}(s,a)\geq\mu/2$, then $1/\sqrt{C^{(n)}_{t,w}(s,a)}\leq 1/\sqrt{\mu/2}=\sqrt{2/\mu}.$ This part of the expectation is therefore bounded by $\sqrt{2/\mu}\cdot P(C^{(n)}_{t,w}(s,a)\geq\mu/2)\leq\sqrt{2/\mu}\leq\sqrt{\frac{2}{w_{n}\mathrm{p}_{min}}}.$
If $C^{(n)}_{t,w}(s,a)<\mu/2$, we can use Chernoff bound to have $P(C^{(n)}_{t,w}(s,a)<(1-1/2)\mu)\leq e^{-\mu/8}\leq e^{-w_{n}P_{\mathrm{min}}/8}$
Thus, the expectation becomes
$\mathbb{E}\left[\frac{1}{\sqrt{C^{(n)}_{t,w}(s,a)}}\right]\leq\frac{\sqrt{2}}{\sqrt{w_{n}P_{\mathrm{min}}}}+e^{-w_{n}P_{\mathrm{min}}/8}\leq\frac{4}{\sqrt{w_{n}P_{\mathrm{min}}}}$ (30)
Therefore, we have
$g_{t,n}(s,a,w)$ $=\mathbb{E}_{P_{n,t},\pi_{n,t}}\left[\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w}(s,a)}}\right]\leq\frac{4H}{\sqrt{w_{n}P_{\mathrm{min}}}}$ (31)
Next, we have
$\sum_{t=1}^{T}\bigg(\frac{4H}{\sqrt{w_{n}P_{\mathrm{min}}}}+\frac{w_{n}V_{n}H}{T}\bigg)=H\bigg(\frac{4T}{\sqrt{w_{n}P_{\mathrm{min}}}}+V_{n}w_{n}\bigg).$ (32)
Then, by substituting $w_{n}=(T/V_{n})^{2/3}$, we get
$H\bigg(\frac{4T}{\sqrt{w_{n}P_{min}}}+w_{n}V_{n}\bigg)=4HT^{2/3}V_{n}^{1/3}P_{\mathrm{min}}^{-1/2}+HT^{2/3}V_{n}^{1/3}=\tilde{O}\bigg(T^{2/3}V_{n}^{1/3}\bigg),$ (33)
where $H$ and $P_{\mathrm{min}}$ are absorbed in big-O-notation because $H$ is constant number time-slots in every episode, $P_{\mathrm{min}}$ depends on the number of time-slots $H$ and the initial state in any episode.
By substituting $V_{n}=\tilde{V}_{n}+2T/J_{n}$ in the above, we obtain Theorem 2.
A.4 Proof of Theorem 3
We first decompose $\mathrm{Term2}$. By adding and subtracting
$\sum_{n=1}^{N}U(\pi_{n,t},\tilde{P}_{n,t},\lambda)~\text{and}~R_{t}\left(\pi_{t},(\tilde{P}_{n,t})_{n=1}^{N}\right),$
we can express $\mathrm{Term2}$ as follows:
$\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}U\left(\pi_{n,t},P_{n,t},\lambda\right)-R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right)\bigg)$
$=h(N)T+\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}U(\pi_{n,t},P_{n,t},\lambda)-\sum_{n=1}^{N}U(\pi_{n,t},\tilde{P}_{n,t},\lambda)$
$+R_{t}\left(\pi_{t},(\tilde{P}_{n,t})_{n=1}^{N}\right)-R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right).$ (34)
When the confidence bound holds,
$U\left(\pi_{n,t},P_{n,t},\lambda\right)-U\left(\pi_{n,t},\tilde{P}_{n,t},\lambda\right)\leq 0.$ (35)
This is because $\pi_{n,t}$ is the optimal solution of the Lagrangian problem and $\tilde{P}_{n,t}$ achieves the highest Lagrangian objective value.
Similar to equation A.2(a), by using [30, Theorem 6.4], we can have
$R_{t}\left(\pi_{t},(\tilde{P}_{n,t})_{n=1}^{N}\right)-R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right)$
$=$ $\mathbb{E}_{P_{t},\pi_{t}}\bigg[\sum_{n=1}^{N}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}}\gamma^{h-1}(\tilde{P}_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h})-P_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h}))V_{n}(s^{\prime};\pi_{t},\tilde{P}_{n,t})\bigg]$
$\leq$ $\mathbb{E}_{P_{t},\pi_{t}}\bigg[\sum_{n=1}^{N}\sum_{(s,a)\mathcal{S}\times\mathcal{A}}\alpha_{t}^{(n)}(s,a)d_{t}^{(n)}(s,a)\bigg]V_{max}$ (36)
which is similar to the last inequality of equation A.2. Hence, similar to Theorem 1 and Theorem 2, we can have
$R_{t}\left(\pi_{t},(\tilde{P}_{n,t})_{n=1}^{N}\right)-R_{t}\left(\pi_{t},(P_{n,t})_{n=1}^{N}\right)\leq\tilde{O}(T^{2/3}(\tilde{V}_{n}+2T/J)^{1/3})+\tilde{O}(\sqrt{TJ_{n}})$ (37)
Therefore, we have
$\mathrm{Reg(T)}$ $\leq\mathrm{Term1}+\mathrm{Term2}$
$\leq\sum_{n=1}^{N}\tilde{O}(2T^{2/3}(\tilde{V}_{n}+2T/J)^{1/3})+\sum_{n=1}^{N}\tilde{O}(2\sqrt{TJ_{n}})+h(N)T.$
Next, we can substitute $J_{n}=O(T^{3/5})$ and obtain
$\tilde{O}(2T^{2/3}(\tilde{V}_{n}+2T/J)^{1/3})+\tilde{O}(2\sqrt{TJ_{n}})$
$\leq\tilde{O}(2T^{2/3}\tilde{V}_{n}^{1/3})+\tilde{O}(2^{4/3}T^{4/5})=\tilde{O}(T^{2/3}\tilde{V}_{n}^{1/3}+T^{4/5})$ (38)
Next, by using $\tilde{V}=\max_{n}\tilde{V}_{n}$, we have $\sum_{n=1}^{N}\tilde{O}(T^{2/3}\tilde{V}_{n}^{1/3}+T^{4/5})=\tilde{O}(T^{2/3}\tilde{V}^{1/3}+T^{4/5})$. This concludes the proof.
A.5 Experimental Setup
@@
Firstly, we discuss how we compute the regret in our numerical studies. Because it is not possible to get the optimal policy of RMAB even if everything is known exactly, we use the best Whittle index policy to demonstrate the regret in Table 1 and Fig. 1. Given a policy, we evaluate the regret of the policy at each episode $t$ by subtracting the total discounted sum reward of all arms obtained by the policy from the total discounted sum reward of all arms received by a Whittle index policy with access to the true transition probabilities. In our simulation, we have used MATLAB. In simulation results of Table 1 and Fig. 1, discount factor of $\gamma=0.99$ and time-slots $H=50$ were considered. Moreover, we considered $V_{n}=V$ for all $n$.
Now, we discuss how we model One Dimensional Bandits and Wireless Scheduling.
One Dimensional Bandits: We consider a modified version of the one dimensional RMAB problem studied in [45, 39]. Each arm $n$ is a Markov process with $K$ states, numbered as $0,1,\ldots,K-1$. For our simulations, we set $K=10$. The reward of an arm increases linearly with the current state, i.e. $r(s,a)=s$. If the arm is activated, then it can evolve from state $s$ to $\min\{s+1,K-1\}$ with probability $q_{n}(t)$ or remain in the same state $s$ with probability $1-q_{n}(t)$. If the arm is not activated, then it evolves from state $s$ to $\max\{s-1,0\}$ with probability $p_{n}(t)$ or remain in the same state $s$ with probability $1-p_{n}(t)$. One-dimensional MDPs of this form are often used in health monitoring and machine monitoring applications [46, 47]. In our simulation, we consider - (i) $V_{n}=35$, (ii) $p_{n}(t)$ changes to $\min\{p_{n}(t-1)+\frac{V_{n}}{4T},1\}$ with probability $0.5$, or it changes to $\max\{p_{n}(t-1)-\frac{V_{n}}{4T},0\}$ with probability 0.5 and (iii) $q_{n}(t)$ changes to $\min\{q_{n}(t-1)+\frac{V_{n}}{4T},1\}$ with probability $0.5$, or it changes to $\max\{q_{n}(t-1)-\frac{V_{n}}{4T},0\}$ with probability 0.5.
Wireless Scheduling: We consider a wireless scheduling problem, where $M$ out of $N$ sources can send their observation to a receiver side over an unreliable channel at every time slot $h\in[H]$ of episode $t\in[T]$. Due to channel unreliability, the observation may not be delivered. The goal of the receiver is to estimate the current signal values of all $N$ sources based the information delivered from the sources. The reward for accurate timely estimation can be modeled as the mutual information between the estimated signal and the actual signal. [48] showed that the mutual information can be determined by using a decreasing function $(\mathrm{log}_{2}(1-\sigma_{n}^{2s_{n,h,t}}))/2$ of Age of Information (AoI) for zero-mean i.i.d. Gaussian random variables with variance $\sigma^{2}_{n}$, where AoI $s_{n,h,t}$ of source $n$ is the time difference between current time $h$ and the generation time of the most recently delivered signal. The AoI value of a source $n$ increases by $1$ if the source $n$ is not scheduled. If the source $n$ is scheduled, the AoI value drops to $1$ with probability $q_{n}(t)$ (successful delivery) or increases by $1$ with probability $1-q_{n}(t)$ (unsuccessful delivery). The parameter $q_{n}(t)$ measures the reliability of channel $n$ at time $t$. In our experiment, we assume that $q_{n}(t)$ is unknown and non-stationary for half of the sources, whereas it is is unknown but stationary for the remaining half. For non-stationary arms, the variance of signal values $\sigma_{n}^{2}=0.9$ is used and the probability of successful transmission $q_{n}(t)$ changes to $\min\{q_{n}(t-1)+\frac{V_{n}}{2T},1\}$ with probability $0.6$, or it changes to $\max\{q_{n}(t-1)-\frac{V_{n}}{2T},0\}$ with probability 0.4; the initial value of $q_{n}(t)=0.1$ is used. For the other half, $q_{n}(t)=1$ is unknown but stationary and $\sigma_{n}^{2}=0.5$.
@@
How we can simplify the maximization problem equation 12: In One Dimensional Bandit applications, we can show that $Q_{n,\lambda,t}(s,0)$ is a decreasing function of $p_{n}(t)$, and $Q_{n,\lambda,t}(s,1)$ is an increasing function of $q_{n}(t)$. This is because the reward $r(s,a)$ is an increasing function of state $s$, which results in $V_{n,\lambda,t}(s)$ being an increasing function of state $s$. Using this structure, we can get closed-form solution of the maximization problem equation 12. For our wireless scheduling problem, the value function $V_{n,\lambda,t}(s)$ is an increasing function of $q_{n}(t)$ because the reward function $r(s,a)$ is a decreasing function of the AoI state $s$. This structure allows us to obtain a closed-form solution for the maximization problem equation 12. Furthermore, a closed-form expression for the Whittle index exists [4]. Though the closed-form expression is obtained for average cost problem, we can use the Whittle index expression from [4] when the discount factor $\gamma\rightarrow 1$. In our experiment, as we have used $\gamma=0.99$, the approximation does not affect the performance of our policy. Consequently, we can determine the Whittle index for our problem using a low-complexity method that exploits these closed-form expressions.
input : $\beta\in(0,1],J=O(T^{3/5})$ (we omit subscript $n$ from $J_{n}$ and $V_{n}$)
1 Initialize $V(i)=V_{max}-(i-1)V_{max}/J$
2 Initialize $w_{i}=1$ and $\hat{X}_{i}=0$
3 for every episode $t=1,2,\ldots,T$ do
4  Set $p_{i}(t)=(1-\beta)\frac{w_{i}}{\sum_{i=1}^{J}w_{i}}+\frac{\beta}{J}$
5   Select $i_{t}\in\{1,\ldots,J\}$ randomly according to probability $p_{1}(t),\ldots,p_{J}(t)$, respectively
6   Select $V(i_{t})$ at episode $t$ and Observe Reward $R_{t}(\pi_{t},(P_{n,t})_{n=1}^{N})$
7   Normalize reward: $X_{t}=R_{t}(\pi_{t},(P_{n,t})_{n=1}^{N})/Nr_{max}H,$ where $r_{max}=\max_{s,a,n}r_{n}(s,a)$
8   Update $\hat{X}_{i_{t}}\leftarrow X_{t}/p_{i_{t}}(t)$
9   Update $w_{i}\leftarrow w_{i}\exp(\beta\hat{X}_{i}/J)$
Algorithm 2 BoB algorithm for choosing $V$